Selected information follow for Bush Company:

2019 2018
Sales revenue $700,000 $520,000
Cost of goods sold 407,700 310,000
Interest expense 22,000 14,000
Income tax expense 6,500 5,100
Net income 30,000 20,300
Cash flow from operating activities 29,500 26,500
Capital expenditures 42,000 25,000
Accounts receivable (net), December 31 182,000 128,000
Inventory, December 31 225,000 180,000
Stockholders’ equity, December 31 205,000 165,000
Total assets, December 31 460,000 350,000


Required
Calculate the following ratios for 2019. The 2018 results are given for comparative purposes.

Round answers to one decimal place. Use 365 days in a year.

2018 2019
1. Gross profit percentage 40.4% Answer %
2. Return on assets 6.5% Answer %
3. Return on sales 3.9% Answer %
4. Return on common stockholders’ equity
(no preferred stock was outstanding) 14.2% Answer %
5. Accounts receivable turnover 4.8 Answer
6. Average collection period 76.5 days Answer days
7. Inventory turnover 2.1 Answer
8. Times-interest-earned ratio 2.8 Answer
9. Operating-cash-flow-to-capital-expenditures ratio 1.1 Answer
Answer #1

First 4 ratios are being calculated here:

1. Gross Profit percentage ratio = Gross Profit / Sales * 100

For 2019:

Sales (given) = $700000

Gross Profit = Sales - Cost of goods sold

Gross Profit = $700000 - $407700 = $292300

Putting the values in the above gross profit percentage formula, we get,

Gross Profit ratio = $292300 / $700000 * 100 = 41.8%

2. Return on assets (ROA) = Net income / Average total assets * 100

where, Average total assets = Beginning assets + Ending assets / 2

So, Average assets = ($350000 + $460000) / 2 = $810000 / 2 = $405000

Net income = $30000

Now, putting these values in the above ROA formula, we get,

Return on assets (ROA) = $30000 / $405000 * 100 = 7.4%

3. Return on sales  = Net income / Sales * 100

Net income = $30000, Sales = $700000

Now, putting these values in the above formula, we get,

Return on sales = $30000 / $700000 * 100 = 4.3%

4. Return on common stockholder's equity (ROE) = Net income / Average stockholder's equity * 100

where, Average stockholder's  equity = Beginning stockholder's equity + Ending stockholder's equity / 2

For 2019:

So, Average stockholder's equity = ($165000 + $205000) / 2 = $370000 / 2 = $185000

Net income = $30000

Now, putting these values in the above ROE formula, we get,

Return on equity (ROE) = $30000 / $185000 * 100 = 16.2%